A custom car built for pretentious people has two brass horns that are supposed to produce the same frequency but actually emit 269.4 and 270.7 Hz. What beat frequency is produced? Hz
Answer #1

beat frequency = f2-f1 = 270.7 -269.4 = 1.3 Hz